OpenStudy (anonymous):

Candice has a dog walking business. It takes her 40 minutes to complete her route with 2 dogs. For each additional dog walked at the same time, it takes her 5 more minutes to complete her route. If x represents the total number of dogs she walks at the same time, which of the following equations can be used to find out how long it takes Candice to complete her route? y = 5(x - 2) + 40 y = 40x + 5 y = 5x + 40 y = 40(x - 2) + 5

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k

OpenStudy (danjs):

A

OpenStudy (anonymous):

Yunika downloaded two applications for $2.99 each to her cell phone. The remaining applications that Yunika purchased cost $1.99 each. If she downloaded x total applications, which of the following equations can be used to find how much Yunika spent on cell phone application downloads? y = 1.99x + 5.98 y = 5.98x + 1.99 y = 5.98(x - 2) + 1.99 y = 1.99(x - 2) + 5.98

OpenStudy (danjs):

D

OpenStudy (anonymous):

Several years ago, Dave bought an antique clock. Since then, it has increased in value. If x represents the number of years he has owned the clock, and y represents the value of the clock (in dollars), which of the following situations is represented by the graph below? Dave bought the clock for $300. Each year the value of the clock increased by $200. Dave bought the clock for $300. Each year the value of the clock increased by $100. Dave bought the clock for $300. Each year the value of the clock increased by $125. Dave bought the clock for $300. Each year the value of the clock increased by $50.

OpenStudy (anonymous):

https://app34.studyisland.com/pics/210386pic2b.png

OpenStudy (danjs):

wait

OpenStudy (danjs):

its D

OpenStudy (anonymous):

A study was done to investigate the population of polar bears over time. The correlating linear model is shown below, where x represents the number of years after 1955, and y represents the number of polar bears. Interpret the y-intercept. y = 5,000 + 500x There were 10,000 polar bears in 1955. There were 1,000 polar bears in 1955. There were 500 polar bears in 1955. There were 5,000 polar bears in 1955.

OpenStudy (danjs):

D

OpenStudy (anonymous):

A company developed the following linear model to help budget annual expenses over time. In this model, x represents the number of years after 2009, and y represents the expense budget (in dollars). Interpret the y-intercept. y = 189,785 + 37,896x The company's expense budget for 2009 was $37,896. The company's expense budget for 2009 was $189,785. The company's expense budget for 2008 was $37,896. The company's expense budget for 2008 was $189,785.

OpenStudy (danjs):

how many questions?

OpenStudy (anonymous):

Around 15?why?

OpenStudy (danjs):

k

OpenStudy (danjs):

The company's expense budget for 2009 was $189,785.

OpenStudy (anonymous):

Unless you can help me with Linear Relationships too,.LOL.Whats the answer to that one

OpenStudy (danjs):

these are linear relationships..lol

OpenStudy (anonymous):

Oh im meant System of equations.xD

OpenStudy (danjs):

ok

OpenStudy (anonymous):

Whats the annswer to the last question

OpenStudy (danjs):

The company's expense budget for 2009 was $189,785.

OpenStudy (anonymous):

Sami consulted a dietitian regarding weight loss and was given a nutrition plan to follow. The dietitian informed Sami, if she follows the plan, her results should be in alignment with the following linear model. In this model, x represents the number of weeks on the plan, and y represents Sami's weight (in pounds). Interpret the y-intercept. y = 342 - 2.2x The y-intercept represents the amount of weight Sami should lose each week. The y-intercept represents Sami's goal weight upon completion of the plan. The y-intercept represents the amount of weight Sami should lose in total. The y-intercept represents Sami's weight upon starting the plan.

OpenStudy (danjs):

D

OpenStudy (anonymous):

Stephanie planted a home garden. She planted 3 rows of tomatoes with 4 tomato plants in each row. She also planted squash in rows of 6 plants each. If x represents the number of rows of squash she planted, which of the following equations can be used to find the total number of plants Stephanie planted in her home garden? y = 6x + 7 y = 12x + 6 y = 7x + 6 y = 6x + 12

OpenStudy (danjs):

D

OpenStudy (anonymous):

For delivery orders, customers must order at least two dozen donuts. The first two dozen donuts cost $4.99 each. Any additional donuts ordered cost $3.99 per dozen. If x represents the number of additional dozens of donuts, which of the following equations can be used to find the cost of the order, not including tax or a delivery charge? y = 4.99x + 3.99 y = 3.99x + 9.98 y = 9.98x + 3.99 y = 3.99x + 4.99

OpenStudy (danjs):

i think the last 10 were all D. hah

OpenStudy (anonymous):

ikr the maker must love D's

OpenStudy (danjs):

haha

OpenStudy (anonymous):

For delivery orders, customers must order at least two dozen donuts. The first two dozen donuts cost $4.99 each. Any additional donuts ordered cost $3.99 per dozen. If x represents the number of additional dozens of donuts, which of the following equations can be used to find the cost of the order, not including tax or a delivery charge? y = 4.99x + 3.99 y = 3.99x + 9.98 y = 9.98x + 3.99 y = 3.99x + 4.99

OpenStudy (danjs):

ah man, it is B this time

OpenStudy (danjs):

any others?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

A study was done to investigate the relationship between outdoor temperature and the amount of fluids an outdoor athlete drinks per day. The correlating linear model is shown below, where x represents the number of degrees over 80°F, and y represents the amount of fluids drank (in liters). Interpret the y-intercept. y = 3.85 + 0.55x An athlete drinks 0.55 L per day when the temperature is 80°F. An athlete drinks 7 L per day when the temperature is 80°F. An athlete drinks 3.85 L per day when the temperature is 80°F. An athlete drinks 4.4 L per day when the temperature is 80°F.

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

C

OpenStudy (anonymous):

Jarrod helped his teachers pack textbooks into boxes for the summer. He packed 6 boxes of science textbooks with 18 books per box. He also packed math textbooks with 10 books per box. If x represents the total number of boxes he packed, which of the following equations can be used to find the total number of books that Jarrod packed into boxes? y = 10x + 108 y = 108x + 10 y = 108(x - 6) + 10 y = 10(x - 6) + 108

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

um

OpenStudy (anonymous):

What

OpenStudy (danjs):

D

OpenStudy (anonymous):

Geoff purchased an annual golf pass for a municipal golf course in his town. He pays a flat fee for the annual golf pass and then each round he plays he must pay the additional cost for a golf cart. A linear model of this situation contains the values (23, 1,152) and (38, 1,317), where x represents the number of times he plays each year, and y equals the total amount he spends on golf in one year. What is the flat fee for the annual golf pass? $1,064 $899 $11 $910

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

899

OpenStudy (anonymous):

Serenity Spa charges $79 for a one-hour massage. They charge $0.80 for each additional minute over an hour. If x represents the number of minutes over one hour, which of the following equations represents the cost of a massage? y = 79x + 0.8 y = -0.8x + 79 y = 0.8x + 79 y = 79.8x

OpenStudy (danjs):

C

OpenStudy (anonymous):

Carrie's Catering is having a special on their box lunches. The minimum purchase is 10 boxes for $70. Each additional box is $5. If x represents the number of boxes over the minimum, which of the following equations represents the cost of box lunches during this special? y = 70 - 5x y = 5x + 70 y = 70x + 5 y = 70x - 5

OpenStudy (danjs):

B

OpenStudy (anonymous):

Georgia filled her kitchen sink up with water so that she could do the dishes. When she was done with the dishes, she pulled out the drain stopper so the water could begin to drain out of the sink. A linear model of this situation contains the values (2, 7.6) and (4, 7.2), where x represents the number of seconds, and y represents the water level in the sink, in inches. What is the rate of change in this linear model? -0.4 of an inch per second 0.2 of an inch per second -8 inches per second -0.2 of an inch per second

OpenStudy (danjs):

D

OpenStudy (anonymous):

The graph below shows the cost of a taxi in New York City. What is the rate of change of the graph? $1.00 per mile $4.00 per mile $2.00 per mile $2.50 per mile

OpenStudy (anonymous):

https://app34.studyisland.com/pics/159235q1a.png

OpenStudy (danjs):

C

OpenStudy (anonymous):

Mathnificent is a math tutoring service. The first year it was open, it had 7 students. If 5 new students enroll each year, which of the following equations represents how many students will be enrolled in x years? y = 7x + 5 y = 5x - 7 y = 7x - 5 y = 5x + 7

OpenStudy (danjs):

D

OpenStudy (anonymous):

Gus has a lunch food budget of $180 per month. He spends an average of $5 per day on lunch. Which of the following equations represents how much money is left in his lunch budget after x days of buying lunch? y = 180 - 5x y = 5x - 180 y = 180x - 5 y = 180 + 5x

OpenStudy (danjs):

A

OpenStudy (anonymous):

Amanda borrowed money from her dad. She has started to pay back her loan, without interest. The amount of the loan she has remaining after each month is represented on the graph below. What is the initial value of the data in the graph? $1,875 $2,000 $125 $750

OpenStudy (anonymous):

https://app34.studyisland.com/pics/159235q11.png

OpenStudy (danjs):

B

OpenStudy (anonymous):

At the beginning of the summer, Jenna had saved $224. To earn money over the summer, she is working part time at the local community center. She will earn $25 per day of work. If x represents the number of days she works, which of the following equations represents how much money she will have, including her savings? y = 25x + 224 y = x + 25 y = 224x + 25 y = 224 - 25x

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

A